OpenStudy (anonymous):

Solve the system of equations: -2x-y=0 3x-y=10

OpenStudy (luigi0210):

Solve for y in the top equation

OpenStudy (anonymous):

x = 2 & y = -4.

OpenStudy (anonymous):

I solved using elimination, but there are 3 other ways.

OpenStudy (anonymous):

Solve the following system: {-2 x-y = 0 | (equation 1) 3 x-y = 10 | (equation 2) Swap equation 1 with equation 2: {3 x-y = 10 | (equation 1) -(2 x)-y = 0 | (equation 2) Add 2/3 × (equation 1) to equation 2: {3 x-y = 10 | (equation 1) 0 x-(5 y)/3 = 20/3 | (equation 2) Multiply equation 2 by 3/5: {3 x-y = 10 | (equation 1) 0 x-y = 4 | (equation 2) Multiply equation 2 by -1: {3 x-y = 10 | (equation 1) 0 x+y = -4 | (equation 2) Add equation 2 to equation 1: {3 x+0 y = 6 | (equation 1) 0 x+y = -4 | (equation 2) Divide equation 1 by 3: {x+0 y = 2 | (equation 1) 0 x+y = -4 | (equation 2) {x = 2 y = -4

OpenStudy (luigi0210):

Haha smartypants

OpenStudy (anonymous):

I'm decent lml. (;

OpenStudy (anonymous):

That's exactly what I got. Thanks so much!